OpenStudy (anonymous):

Let f(x)=-4+7 and g(x)=10x-6 , find f(g(x)). A. -40x+24 B. -40x+31 C. -40x+64 D. -40x+70 *NEED THIS ANSWER ASAP , PLUS HELP*

OpenStudy (anonymous):

\[f(g(x))= f(10x-6)= -4(10x-6) + 7 = -40x+31\] correct the question \[f(x) = -4x+7\]